OpenStudy (anonymous):

whats is the value of theta , if cos theta = -1/2

OpenStudy (jack1):

cos theta = -1/2 so theta = cos^-1 (-1/2)

OpenStudy (anonymous):

let me use x instead of theta cosx=-1/2 cosx=-cos(pi/3) cosx = cos(pi- pi/3)=cos2(pi/3)

OpenStudy (anonymous):

i.e cos x =cos 2(pi/3) or x= 2(pi/3)